OpenStudy (anonymous):

A jar contains 2 orange balls, 5 black balls, and 3 pink balls. Event A = drawing a black ball on the first draw Event B = drawing a pink ball on the second draw If two balls are drawn from the jar, one after the other without replacement, what is P(A and B) expressed in simplest form?

OpenStudy (anonymous):

Event A= draw a black ball on the first draw. = 5/ (2+5+3) = 0.5 = 50% Event B= draw a pink ball on the second draw. = P(O, P) +P(B, P) + P(P,P) = probability of drawing orange first then pick + probability of drawing black then pink + probability of drawing pink then pink again = (2/10)(3/9) + (5/10)(3/9) + (3/10)(2/9) = 0.3 = 33% P (event A and B) = P (drawing a black then a pink) = (5/10)(3/9)= 16.6%